OpenStudy (anonymous):

0.2(60) + 0.05x = 0.10(60 + x)

OpenStudy (anonymous):

First resolve the brackets to get 12 + 0.05x = 6 + 0.10x7 Now separate the x terms on one side and constants on the other 0.05x - 0.10x = 6 - 12 - 0.05x = - 6 x = -6/-0.05 x = 120